OpenStudy (anonymous):

(just need help explaining what formula I have to do to get my answer.) Glenn invested $22,000 at 7% interest compounded annually. How much interest will Glenn earn in 3 years?

OpenStudy (astrophysics):

\[\huge A = P(1+\frac{ r }{ n })^{nt}\] compound interest formula should suffice. P = Principle amount r = Annual rate of interest t = Number of years A = Amount of money accumulated after n years n = Number of times the interest is compounded per year

OpenStudy (anonymous):

$$\huge A = P \left( 1+\frac{ r }{ n } \right)^{nt}$$

OpenStudy (anonymous):

sorry confused am I supposed to use the 1 in every formula so that I can get my answer?

OpenStudy (astrophysics):

Yes!

OpenStudy (anonymous):

how do I find n?

OpenStudy (astrophysics):

`Glenn invested $22,000 at 7% interest compounded annually` what does annually mean

OpenStudy (anonymous):

every year so 22000 x 3 would be n?

OpenStudy (astrophysics):

No the interest is compounded annually so n = 1

OpenStudy (anonymous):

make sure ur writing ur interest in decimal form

OpenStudy (anonymous):

im doing the formula but I must be doing something wrong because none of the answers I get are right

OpenStudy (astrophysics):

Let me ask you, which variable are we solving for

OpenStudy (anonymous):

um sorry I don't really know variables but I'm thinking 3 is the variable ??

OpenStudy (astrophysics):

Are we solving for A, p, n, r, t?

OpenStudy (anonymous):

a? idk we are solving what 22,000 at 7% would be in 3 years

OpenStudy (anonymous):

\[A=22000(1+\frac{ 0.07 }{ 1 })^3\]

OpenStudy (astrophysics):

Yes @mavckenzie, ayesh has plugged in all the values for you please see if you understand it.